You manage a manufacturing Production department making a family of similar parts. You must make decisions about equipment selection, production cost and sequence of production jobs
Part material cost is $ 12 for each scrapped

Options

MACHINE LABOR COST PRODUTION RATE PART PART
(Per hr) (Units per hr) scrap rate scrap cost
1.Manual $20 10 L out of 32 $10
2.CNC with manual load $18 12 L out of 80 $5
3.CNC with auto lo9ad $15 15 L out of 60 $12
4.Dedicated automation $12 60 L out of 120 $25
Answer #1

1. This is a simple problem and the decision should be to go ahead with Dedicated Automation. Assuming the fact that the demand is there to sell all of the produced units, with reduced machine cost compared to manual labor, it helps in producing 6 times more units. Even though the scrap cost is $15 more than the Manual one, the scrap rate is the same L units but for 120 units of production.

Lets take the example of 120 units of production. Assuming L is 1 here as it is the same throughout all methods

Cost for Manual - $ 20 * 12 hours pf production + $10 scrap cost * (120/32) = 240 + 37.5 = $ 277.5

Cost for CNC manual = $ 18 * 10 hours of production + $ 5 scrap cost *(120/80) = 180+7.5 = $ 187.5

Cost for CNC Auto = $15* 8 hours of production + $ 12* (120/60) = 120 + 24 = $144

Cost for Automated = $ 12 * 2 hours of production + $25*(120/120) = 24 + 25 = $ 49

As proved in the above example, Dedicated Automation has the least amount of cost involved and hence is an easy selection